# AutoPrep: Natural Language Question-Aware Data Preparation with a Multi-Agent Framework

### v1

@@ 7. Related Work
Tabular Question Answering. Most of the SOTA solutions for TQA rely on Large Language Models (LLMs) (Brown et al., 2020; Guo et al., 2024), as TQA requires NL understanding and reasoning over tables. The existing solutions can be broadly divided into two categories: end-to-end reasoning (Wei et al., 2022) and code generation (Ye et al., 2023; Yao et al., 2023; Cheng et al., 2023; Zhang et al., 2024). End-to-end reasoning approaches, which prompt an LLM to directly generate answers through next token prediction (Wei et al., 2022), may struggle when processing large tables due to the input token limitation and long-context understanding challenges (Li et al., 2024). On the other hand, code generation approaches leverage LLMs to generate programming code (e.g., Python) that extracts answers from tables, providing better explainability and achieving higher accuracy on real-world TQA tasks (Cheng et al., 2023; Zhang et al., 2024). For example, Dater (Ye et al., 2023) proposes to decompose questions and use code to deconstruct table contexts, enhancing end-to-end table reasoning with decomposed sub-tables and sub-questions. ReAcTable (Zhang et al., 2024) utilizes the ReAct paradigm (Yao et al., 2023) to generate codes which are executed step-by-step to extract intermediate question-related sub-tables for deducing the answer.
However, current LLM-based TQA methods mainly focus on the reasoning process of answer extraction instead of thorough and comprehensive data prep. While Binder (Cheng et al., 2023) proposes integrating SQL with an LLM-based API to incorporate external knowledge, which may partly address augmentation tasks, it does not account for a wide range of data prep tasks, such as normalization and column filtering. In this paper, we highlight the significance of question-aware data prep for TQA which could be a bottleneck restricting the performance. We also point out primary challenges in question-aware TQA data prep and propose an automatic multi-agent framework to solve them.
@@
Traditional Data Preparation. Data Prep techniques have been widely adapted to a wide range of tasks, such as machine learning (ML) and exploratory data analysis (EDA), as they turn raw data into a format that is ready for downstream tasks. A complete data preparation process usually involves a series of steps, such as data transformation, cleaning, and feature engineering. Auto-Weka (Thornton et al., 2013) leverages Bayesian optimization to identify optimized data prep operation plan. Other approaches, such as Auto-Sklearn (Feurer et al., 2020; Hutter et al., 2019) and TensorOBOE (Yang et al., 2020), apply meta-learning techniques to discover promising data prep configurations. Alpine Meadow (Shang et al., 2019) introduces an exploration-exploitation strategy, while TPOT (Olson and Moore, 2016) uses a tree-based representation of data prep and genetic programming optimization techniques. Several studies (Berti-Équille, 2019; El et al., 2020; Heffetz et al., 2020) have explored the use of reinforcement learning to generate plans autonomously. For human-in-the-loop data prep, systems like DORIAN (Redyuk et al., 2022), Auto-Suggest (Yan and He, 2020), and Auto-Pipeline (Yang et al., 2021) incorporate recommendation-based methods to assist users. HAIPipe (Chen et al., 2023) integrates both human-orchestrated and automatically generated data prep operations.
@@
LLMs-based Multi-Agent Framework. A multi-agent LLM framework refers to a well-designed hierarchical structure consisting of multiple LLM-based agents and scheduling algorithms (Wang et al., 2024a). Compared with single-agent methods based on prompting techniques, such structures are better suited for handling complex tasks like software development, issue resolution, and code generation (Hong et al., 2023; Qian et al., 2023; Chen et al., 2024; Tao et al., 2024; Ishibashi and Nishimura, 2024; Islam et al., 2024). AutoTQA (Zhu et al., 2024) propose a multi-agent framework supporting Autonomous Tabular Question Answering, which decomposes complex TQA problems step-by-step and design agents to identify and correct shortcomings in the problem-solving process.
While AutoTQA improves table analysisx, this paper focuses on the performance bottleneck caused by data prep issues in TQA. Our proposed framework, AutoPrep, addresses question-aware data prep for TQA and can be integrated as a plugin into current TQA approaches to further improve the overall performance.
@@ 8. Conclusion
In this paper, we have introduced AutoPrep, an LLM-based multi-agent framework designed to support automatic data prep for TQA tasks. AutoPrep consists of three stages: (1) the Planning stage, which suggests high-level data prep operations; (2) the Programming stage, which generates low-level implementations for each high-level operation; and (3) the Executing stage, which runs the Python code and iteratively debugs. We also propose a Chain-of-Clauses method to generate high-quality logical plans and a Tool-augmented method for effective physical plan generation. Additionally, we introduce an effective method to ensure successful execution of the physical plan. Extensive experiments on WikiTQ and TabFact demonstrate that: (1) AutoPrep achieves state-of-the-art performance in data prep for TQA, and (2) integrating AutoPrep’s data prep improves the overall performance of TQA methods.

### v2

@@ 7. Related Work
Tabular Question Answering. Most of the SOTA solutions for TQA rely on LLMs (Brown et al., 2020; Guo et al., 2024), as TQA requires NL understanding and reasoning over tables. Specifically, direct Prompting approaches (Chen, 2022; Wei et al., 2022), which prompt an LLM to directly generate answers through next token prediction, may struggle when processing large tables due to the input token limitation and long-context understanding challenges (Li et al., 2024). Code Generation approaches (Rajkumar et al., 2022) leverage LLMs to generate programming code (e.g., Python) that extracts answers from tables, providing better explainability and adaptability.
Previous TQA methods, like Dater (Ye et al., 2023), Binder (Cheng et al., 2023), CoTable (Wang et al., 2024b) and ReAcTable (Zhang et al., 2024), also consider data prep in their question answering process. Specifically, Dater (Ye et al., 2023) prompts LLMs to select relevant columns related to answering the question, targeted at solving filtering tasks. Binder (Cheng et al., 2023) proposes to integrate SQL with an LLM-based API to incorporate external knowledge, which may partly address augmentation tasks. Similarly, CoTable (Wang et al., 2024b) addresses the filtering tasks and augmentation tasks by designing operators which are implemented by LLM completion. ReAcTable (Zhang et al., 2024) uses few-shot demonstrations to instruct LLMs to generate python code or SQL to address the augmentation and filtering tasks.
However, previous TQA methods have the following limitations when handling data prep. First, all these methods do not address the cell-level normalization tasks, which account for the most significant error types, as indicated in Figure 1. Second, all these methods utilize a single LLM agent for both data prep and answer reasoning. Given that data prep is a complex challenge, these one-size-fits-all solutions still fail to achieve satisfactory performance as demonstrated in our experiments.
@@
Traditional Data Preparation. Data prep techniques have been widely adopted in various tasks. Auto-Weka (Thornton et al., 2013) leverages Bayesian optimization to identify data prep operations. Auto-Sklearn (Feurer et al., 2020; Hutter et al., 2019) and TensorOBOE (Yang et al., 2020) apply meta-learning to discover promising operations. Alpine Meadow (Shang et al., 2019) introduces an exploration-exploitation strategy, while TPOT (Olson and Moore, 2016) uses a tree-based representation of data prep and genetic programming optimization techniques. Several studies (Berti-Équille, 2019; El et al., 2020; Heffetz et al., 2020) explore reinforcement learning techniques. HAIPipe (Chen et al., 2023) integrates both human-orchestrated and automatically generated data prep operations.
@@
LLMs-based Multi-Agent Framework. A multi-agent LLM framework refers to a well-designed hierarchical structure consisting of multiple LLM-based agents and scheduling algorithms (Wang et al., 2024a). Compared with single-agent methods based on prompting techniques, such structures are better suited for handling complex tasks like software development, issue resolution, and code generation (Hong et al., 2023; Qian et al., 2023; Chen et al., 2024; Tao et al., 2024; Ishibashi and Nishimura, 2024; Islam et al., 2024). AutoTQA (Zhu et al., 2024) propose a multi-agent framework supporting Tabular Question Answering.
While AutoTQA improves table analysis, this paper focuses on the performance bottleneck caused by data prep issues in TQA. Our proposed framework, AutoPrep, addresses question-aware data prep for TQA and can be integrated as a plugin into current TQA approaches to further improve the overall performance.
@@ 8. Conclusion
In this paper, we have introduced AutoPrep, an LLM-based multi-agent framework designed to support automatic data prep for TQA tasks. AutoPrep consists of three stages: (1) the Planning stage, which suggests high-level data prep operations; (2) the Programming stage, which generates low-level implementations for each high-level operation; and (3) the Executing stage, which runs the Python code and report error messages. We also propose a Chain-of-Clauses method to generate high-quality logical plans and a Tool-augmented method for effective physical plan generation. Additionally, we introduce an effective method to ensure successful execution of the physical plan. Extensive experiments on WikiTQ, TabFact and TabBench demonstrate that: (1) AutoPrep achieves SOTA performance in data prep for TQA, and (2) integrating AutoPrep’s data prep improves the performance of TQA methods.